OpenStudy (anonymous):

Which is the direct linear variation equation for the relationship? y varies directly with x and y = 12 when x = 4. A. y = x + 8 B. y = 2x + 4 C. y = 3x D. y = x – 8

OpenStudy (anonymous):

@iGreen @NeonStrawsForever @razorghost @rvc @tylermcmullen23

rvc (rvc):

do you know the equation for direct variation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it y=mx+b

OpenStudy (anonymous):

i think it might be C.

rvc (rvc):

\[\huge y~ \alpha~ x\]

OpenStudy (anonymous):

y=a+x?

rvc (rvc):

therefore y=kx where k is a constant

OpenStudy (anonymous):

y varies directly with x and y = 12 when x = 4. (Points: 4) y=kx y/x=k 3=k Option C: y = 3x

OpenStudy (anonymous):

so it would be c then?

OpenStudy (anonymous):

id say (:

rvc (rvc):

@Crumby no direct answers please

OpenStudy (anonymous):

oh i remember this now thank you guys

OpenStudy (anonymous):

i gave a problem for her/he to figure out and the answer is Attached no direct answers at all (;

rvc (rvc):

All the best !!!

rvc (rvc):

you got to explain her

OpenStudy (dmndlife24):

A direct answer - with or without an explanation - is still a violation of the CoC

OpenStudy (anonymous):

yes im just here to help and give an explanation

OpenStudy (anonymous):

did not read that at all been on here for a hour thats it so i honestly don know where im at fault

rvc (rvc):

okay but now you know that giving direct answers is against the rules never do that again :)